At April 30, partners’ capital balances in Blossom Company are G. Donley $55,000, C. Lamar $45,000, and J. Pinkston $17,600. The income sharing ratios are 5 : 4 : 1, respectively. On May 1, the PDLT Company is formed by admitting J. Terrell to the firm as a partner.

Journalize the admission of Terrell under each of the following independent assumptions. (Credit account titles are automatically indented when amount is entered. Do not indent manually. Round answers to 0 decimal places, e.g. 5,275.)
(1) Terrell purchases 50% of Pinkston’s ownership interest by paying Pinkston $15,000 in cash.
(2) Terrell purchases 331/3% of Lamar’s ownership interest by paying Lamar $15,000 in cash.
(3) Terrell invests $62,800 for a 30% ownership interest, and bonuses are given to the old partners.
(4) Terrell invests $44,600 for a 30% ownership interest, which includes a bonus to the new partner.
Answer #1
NO Account Title And Explanation Debit Credit
1) J. Pinkston Capital ( $ 17,600 * 50% ) $ 8,800
                             Terrell Capital $ 8,800
( To Record Terrell purchase 50% of Pinkston's ownership Interest )
2) C.Lamar Capital ($45,000 *33.33% ) $ 14,999
                             Terrell Capital $ 14,999
( To Record Terrell purchase 33.33% of Lamar's ownership Interest )
3) Cash $ 62,800
          Terrell Capital $ 54,120
         G.Donley Capital $ 4,340
         C.Lamar Capital $ 3,472
          J.Pinkston Capital $ 868
( To record Terrell invest $ 62,800 for 30% ownership interest
and bonus are given to the old partners { All explanation in note 2})
4) Cash $ 44,600
G.Donley Capital $ 2,030
C.Lamar Capital $ 1,624
J.Pinkston Capital $ 406
                 Terrell Capital $ 48,660
( To record Terrell invest $ 44,600 for 30% ownership interest,
which includes bonus are given to the New partners { All explanation in note 3})
Notes :
1) Here Transaction 1 and 2 , these are indirectly from the partnership.
( 50% of $ 17,600 , 33.33% of $ 45,000 ) of respective partner's capital
It is inrelevant for amount of received by the partner.
2) Terrell Capital = ($ 55,000+$ 45,000 +$ 17,600 )+ $ 62,800 = $180,400
Terrell Capital = $ 180,400 *30 % = $ 54,120
Bonus to old partner capital = $ 62,800 - $ 54,120 = $ 8,680
G.Donley Capital = $ 8,680 * 5/10 = $ 4,340
C.Lamar Capital = $ 8,680 *4/10 =$ 3,472
J.Pinkston Capital =$ 8,680 *1/10 = $ 868
3) Terrell Capital = ($ 55,000+$ 45,000 +$ 17,600 )+ $ 44,600 = $ 162,200
Terrell Capital = $ 162,200 *30 % = $ 48,660
Bonus to New partner( Terrell capital = $ 48,660 - $ 44,600= $ 4,060
G.Donley Capital = $ 4,060 * 5/10 = $ 2,030
C.Lamar Capital = $ 4,060 *4/10 =$ 1,624
J.Pinkston Capital =$ 4,060 *1/10 = $ 406